OpenStudy (anonymous):

Find the derivative of the function: s(x) = x(x^2 - 1/x)

OpenStudy (anonymous):

mutiply out first and get \[s(x)=x^3-1\]

OpenStudy (riley):

Once you've multiplied through like Satellite was saying then you can find the derivative I believe. s(x) = x^3 -1 s'(x) = 3x^2 So the derivative = 3x^2 And thats what your left with since the -1 goes away. I think I did that correctly . . .